OpenStudy (anonymous):

y=-x-5 and y=x+1 ( ordered pairs) ?

OpenStudy (lgbasallote):

y = -x - 5 and y = x + 1... do you agree that y = y?

OpenStudy (anonymous):

thats the same thing , i dont get it .

OpenStudy (lgbasallote):

so you do agree that y = y therefore -x - 5 = x + 1 right?

OpenStudy (lgbasallote):

because -x - 5 is y...and x+1 is also y

OpenStudy (anonymous):

so if its graphed that will be the ordered pairs ?

OpenStudy (anonymous):

no it needs a X and Y !

OpenStudy (lgbasallote):

we'll get to that

OpenStudy (lgbasallote):

-x - 5 = x + 1 we solve for x first -1 - 5 = x + x -6 = 2x x = -3

OpenStudy (lgbasallote):

now sub this into y = x + 1 y = -3 + 1 y = -2

OpenStudy (lgbasallote):

therefore your ordered pair is (-3,2)

OpenStudy (lgbasallote):

(-3,-2) *

OpenStudy (anonymous):

urgh , wish i was smart , because i just dont get it :( like how to soolve for it. but thanks.